# Determination of galaxy photometric redshifts using Conditional Generative Adversarial Networks (CGANs)

### v1

@@ 4 Conclusions
On this paper, it was presented a new machine-learning technique for photometric redshift estimation using Conditional Generative Adversarial Networks (CGANs), was presented. Propsed CGAN techique was tested using Dark Energy Survey Y1 data and compared with the classic machine-learning technique the Random Forest Regressor.
Results obtained at CGAN approach, exhibit a similar performance for determining the mean photometric redshift as the Random Forest Classifier. Whereas Random Forest Regressor can only provide a the expected redshift for each individual galaxy, the proposed CGAN approach is able to provide a full probability density function. This additional capability of the CGAN, allows for a better control for the impact of thick tails of the distribution, making this type of photometric redshift algorithmic approach more suitable for cosmological probes that are specially sensitive to thick tails at photometric redshifts -such as number-counts magnification-.
In addition to the value added, CGAN demonstrates a superior computing performance over Random Forest Regressor, as modern Deep Learning frameworks -such as PyTorch-, have a strong parallelization optimization for the determination of the weights of the tensors defining the neural networks, whereas algorithms such as the Random Forest Regressor are not parallelization. This, makes the CGAN approach superior from the computing performance point of view, which is specially suitable for large Cosmology surveys.
It is worth remarking that the results of this work do not constitute a full out-of-the-box software for photometric redshift determination, rather than a proof of concept and a feasibility analysis. This algorithm as still room for improvement in terms of calibration sample selection and the inclusion of additional features.
Accuracy metrics obtained by CGAN, exhibit a perfomance ranking with the best photometric redshift trained on similar calibration data. Performance and accuracy metrics obtained by the CGAN approach described here, is strongly limited by the resolution of the band-pass filters used for the photometry and the galaxy sample used. On this dataset, only four band-pass filters $griz$ are available, thus, much better photometric redshift could be obtained by using a higher resolution apparatus, like the 40-band-pass filter PAUCam.
Future work will explore the possibility of including a finer selection by population of galaxies (main-sequence, LRGs, etc.) that can be included as an additional parameter of the CGAN. Additional features to be explored will also include an evaluation of the inclusion of additional photometric features of the galaxies, such as shape parameters.
Acknowledgements

### v2

@@ 4 Conclusions
On this paper, it was presented a new machine-learning technique for photometric redshift estimation using a Conditional Generative Adversarial Network (CGAN). Propsed CGAN techique was tested using Dark Energy Survey Y1 data and compared with a Mixture Density Network (MDN).
Both point-estimation and probability-density quality-metrics indicate CGAN performance is close to the MDN algorithm. Although these metrics clearly demonstrate that MDN performs better than the proposed CGAN, this work constitutes a proof of concept indicating that the use of CGAN with the $f$-divergence formalism is an algorithm that can be used for photometric redshift determination, opening the door to further developments on the use of CGANs at this field.
Future work will explore the possibility of including a finer selection by population of galaxies (main-sequence, LRGs, etc.) that can be included as an additional parameter of the CGAN that shall be computed as a previous step to the photometric redshift CGAN.
Acknowledgments